# On the Sharp Input-Output Analysis of Nonlinear Systems under Adversarial Attacks

### v1

@@ Abstract
This paper is concerned with learning the input-output mapping of general nonlinear dynamical systems. While the existing literature focuses on Gaussian inputs and benign disturbances, we significantly broaden the scope of admissible control inputs and allow correlated, nonzero-mean, adversarial disturbances. With our reformulation as a linear combination of basis functions, we prove that the $\ell_{1}$-norm estimator overcomes the challenges as long as the probability that the system is under adversarial attack at a given time is smaller than a certain threshold. We provide an estimation error bound that decays with the input memory length and prove its optimality by constructing a problem instance that suffers from the same bound under adversarial attacks. Our work provides a sharp input-output analysis for a generic nonlinear and partially observed system under significantly generalized assumptions compared to existing works.
@@ 1 Introduction
Dynamical systems describe how the state of a system evolves over time according to specific laws. Such systems are ubiquitous in scientific and engineering disciplines, including computer networks (Low et al., 2002), deep learning (Meunier et al., 2022), portfolio management (Grinold and Kahn, 2000), biology (Murray, 2007), and optimal control (Dorf and Bishop, 2011). In many practical settings, however, the underlying dynamics are too complex to be explicitly characterized, resulting in models with partially or entirely unknown parameters. Designing controllers or making predictions without first identifying these unknowns can lead to suboptimal or even unsafe outcomes. To address this challenge, the field of system identification focuses on identifying system dynamics from observed input-output data.
There has been extensive research in system identification under various structural and disturbance assumptions (Simchowitz et al., 2018; Faradonbeh et al., 2018; Simchowitz et al., 2019; Jedra and Proutiere, 2020; Sarkar et al., 2021; Oymak and Ozay, 2022; Yalcin et al., 2024; Zhang et al., 2024; Kim and Lavaei, 2024, 2025) While these works provide strong theoretical guarantees and practical algorithms, the majority of them concentrate on linear systems. However, many real-world systems are inherently nonlinear (Grinold and Kahn, 2000; Low et al., 2002; Murray, 2007), which motivates us to develop identification methods that go beyond the linear setting.
@@
$x_{t+1}$ $=f(x_{t},u_{t},w_{t}),\quad y_{t}=g(x_{t},u_{t}),\quad t=0,1,\dots,T-1,$ (1)
@@
where $x_{t}\in\mathbb{R}^{n}$ is the state, $u_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, and $y_{t}\in\mathbb{R}^{r}$ is the observation at time $t$. The set $\mathcal{U}$ consists of admissible control inputs, and $T$ is the time horizon. Under adversarially chosen disturbances $w_{t}\in\mathbb{R}^{d}$, our goal is to identify the input-output mapping of the system (1).
Nonlinear system identification approaches can generally be classified into two broad categories: nonparametric and parametric methods. While nonparametric approaches offer high flexibility (Greblicki and Pawlak, 2008; Ziemann et al., 2022), they often come with substantial computational overhead and limited interpretability. In contrast, parametric methods use prior knowledge or structural assumptions to yield models that are more interpretable, computationally efficient, and analytically tractable (Chen, 1995; Giannakis and Serpedin, 2001).
In this paper, we adopt a parametric approach, modeling the nonlinear system using a linear combination of basis functions. In principle, we reformulate the input-output mapping from (1) as:
$y_{t}=G^{*}\cdot\Phi(u_{t},\dots,u_{t-\tau})+\textit{extra terms},$ (2)
@@
where $M$ is the number of basis functions, $\Phi:(\mathcal{U})^{\tau+1}\to\mathbb{R}^{M}$ is the stack of basis functions, $G^{*}\in\mathbb{R}^{M\times r}$ represents the matrix governing the input-output mapping, $\tau>0$ is input memory length, and the extra terms are functions of only disturbances and the “far” past states. Note that the extra terms are exponentially small with the exponent $\tau$ under stability conditions when all disturbances are zero. We defer the justification for this reformulation to Section 1.1 below and provide a detailed derivation in Section 2.
Using this formulation, the original nonlinear system is now replaced with a linear combination of basis functions, reducing the system identification task to estimating $G^{*}$. However, the partial observability of the nonlinear system, combined with adversarial disturbances $w_{t}$ and the restriction on admissible control inputs $u_{t}$, introduces significant challenges in accurately recovering $G^{*}$. Thus, we pose the following central question:
When and how can we accurately estimate the true $G^{*}$
under non-Gaussian inputs and correlated, nonzero-mean, adversarial disturbances?
@@
Dynamics Available Outputs Gaussian input i.i.d. Disturbance Zero-mean Disturbance Identification Approach
@@
Oymak and Ozay (2022) Linear Partial Yes Yes Yes Parametric
@@
Kim and Lavaei (2025) Linear Partial Yes No No Parametric
@@
3) We establish the estimation error of $\Theta(\rho^{\tau})$ in nonlinear systems using the $\ell_{1}$-norm estimator, where $10<\rho<$ 1 is the contraction factor determined by the function $f$. This result generalizes prior findings in the linear system identification under Gaussian control inputs and adversarial disturbances (Kim and Lavaei, 2025). We further provide a matching lower bound of $\Theta(\rho^{\tau})$ on the estimation error, showing that the presented bound is indeed optimal.
@@
Outline. The paper is organized as follows. In Section 2, we formulate the problem and state the relevant assumptions. In Section 3, we prove that the $\ell_{1}$-norm estimator achieves the optimal estimation error and provides the analysis outline. In Section 4, we present numerical experiments to validate our main results. Finally, concluding remarks are provided in Section 5.
@@
Notation. Let $\mathbb{R}^{n}$ denote the set of $n$-dimensional vectors and $\mathbb{R}^{n\times n}$ denote the set of $n\times n$ matrices. For a matrix $A$, $\|A\|_{F}$ denotes the Frobenius norm of the matrix. For a vector $x$, $\|x\|_{1}$ denotes the $\ell_{1}$-norm and $\|x\|_{2}$ denotes the $\ell_{2}$-norm of the vector. For a scalar $z$, $\text{sgn}(z)=1$ if $z>0$, $\text{sgn}(z)=-1$ if $z<0$, and $\text{sgn}(z)=0$ if $z=0$. For a set $S$, the $k$-fold Cartesian product $S\times S\times\cdots\times S$ (with $k$ factors) is denoted by $(S)^{k}$. Let $\mathbb{S}^{n-1}$ denote the set $\{y\in\mathbb{R}^{n}:\|y\|_{2}=1\}.$ For an event $E$, the indicator function $\mathbb{I}\{E\}$ equals $1$ if $E$ occurs, and 00 otherwise. $\mathbb{P}(E)$ denotes the probability that the event occurs. We use $\Theta(\cdot)$ for the big-$\Theta$ notation. Let $I_{n}$ denote the $n\times n$ identity matrix. The notation $\succeq$ denotes positive semidefiniteness. Let $N(\mu,\Sigma)$ denote the Gaussian distribution with mean $\mu$ and covariance $\Sigma$, and $\text{Unif}[a,b]^{n}$ denote the uniform distribution on the hypercube $[a,b]^{n}\subset\mathbb{R}^{n}$. The notation $\equiv$ means that the two sides are identically equal. Finally, let $\mathbb{E}$ denote the expectation operator.
1.1 Model Justification
Observing that $y_{t}$ in (1) is transferred from the control inputs $u_{t},u_{t-1},\cdots,u_{0}$, disturbances $w_{t-1},w_{t-2},\cdots,w_{0}$, and the initial state $x_{0}$, we can rewrite $y_{t}$ as a nonlinear function of $u_{t},u_{t-1},\cdots,u_{0},w_{t-1},w_{t-2},\cdots,w_{0},x_{0}$. While this formulation captures all information, it suffers from several key limitations. First, the representation is inherently time-varying and becomes increasingly high-dimensional as time progresses. This not only leads to growing storage and computational burdens but also makes estimation statistically unstable due to the curse of dimensionality and variance accumulation over long horizons. Second, the underlying nonlinear function may be highly complex, potentially involving a large number of parameters, which makes accurate estimation challenging with finite data. Third, the function may involve complex interactions between control inputs and disturbances, further complicating the functional form and hindering identifiability. Motivated by these limitations, we reformulate the model in (1) as shown in (2), which effectively addresses these challenges in the following approaches:
For the first challenge, we adopt a finite-memory approximation strategy with length $\tau$, consistent with classical system identification techniques such as Volterra series truncations (Boyd and Chua, 1985) and NARMAX models (Billings, 2013). These methods are grounded in the assumption that the dynamics of a nonlinear system can effectively be represented using a fixed window of past inputs. By relying on mild regularity conditions, this approach can preserve approximation accuracy while ensuring computational tractability (note that the error term due to using a fixed window of the past is exponentially small when all disturbances are zero).
For the second challenge, we approximate the unknown nonlinear function as a linear combination of basis functions, allowing for flexible yet interpretable modeling. Our primary motivation in using linear combination of basis functions to represent nonlinear functions is rooted in function approximation theory, which ensures that a wide class of nonlinear mappings can be approximated with arbitrary precision using a finite set of appropriately chosen basis functions. For instance, Fourier expansions represent highly nonlinear periodic functions as linear combinations of sine and cosine functions. Radial Basis Function (RBF) networks model nonlinear relationships using linear combinations of radially symmetric basis functions centered at specific points (Chen et al., 1991). In control and signal processing, Volterra series use Volterra kernels to generalize linear system models (Boyd and Chua, 1985). Moreover, random feature methods achieve function approximation through linear combinations of basis functions in a randomized feature space. (Rahimi and Recht, 2007).
To address the third challenge, we design the basis functions to depend solely on the control inputs, thereby decoupling the control inputs and disturbances. This separation has proven effective and is widely adopted in various settings. For example, in model-based reinforcement learning (RL), it is common to alternate between system identification and control policy design, where the agent first learns an input-output model of the environment and then uses it to make informed decisions (Moerland et al., 2023). In many cases, it suffices to focus on the input-output relationship—how control actions affect observable outcomes—rather than attempting to recover the full latent state dynamics (Abbeel et al., 2006; Deisenroth and Rasmussen, 2011). This simplification is particularly valuable in high-stakes applications like autonomous driving, where control inputs such as throttle and steering are mapped to observations such as heading direction, position, and velocity (Paden et al., 2016). By isolating the controllable aspects of the system, this formulation improves both estimation accuracy and robustness to disturbances, especially under adversarial settings.
@@ 2 Problem Formulation
In (1), we study a nonlinear dynamical system $x_{t+1}=f(x_{t},u_{t},w_{t})$ and $y_{t}=g(x_{t},u_{t})$, where the state equation is governed by the dynamics $f:\mathbb{R}^{n}\times\mathcal{U}\times\mathbb{R}^{d}\to\mathbb{R}^{n}$ and the observation equation is determined by the measurement model $g:\mathbb{R}^{n}\times\mathcal{U}\to\mathbb{R}^{r}$. We have the discretion to design control inputs $u_{0},u_{1},\dots,u_{T-1}$ and we have access to a single observation trajectory consisting of partial observations $y_{0},y_{1},\dots,y_{T-1}$. We assume the Lipschitz continuity of the measurement model $g$ and the contraction property for the dynamics $f$ to ensure system stability and prevent the explosion of the nonlinear system, which is common in a number of control theory literature (Sontag, 2008; Khalil, 2015; Tsukamoto et al., 2021). The formal assumption on the dynamics is given below.
@@
$g$ is Lipschitz continuous; i.e. there exists $L>0$ such that
@@
for all $x,\tilde{x}\in\mathbb{R}^{n}$, $u,\tilde{u}\in\mathcal{U}$. Moreover, note that for $k\geq 1$, the $k$-fold composition of the dynamics $f$, denoted by $f^{(k)}$, maps $(x_{t-k},u_{t-k},\dots,u_{t-1},w_{t-k},\dots,w_{t-1})$ to $x_{t}$. We assume that $f^{(k)}$ is Lipschitz continuous in its oldest arguments $(x_{t-k},u_{t-k},w_{t-k})$ with constant $C\rho^{k}$ for some $C>0$ and $10<\rho<$ 1, with later inputs and disturbances fixed. In other words, we have
@@
We consider both inputs and disturbances on the system to be sub-Gaussian variables, in which the tail event rarely occurs. For example, any bounded variables automatically satisfy this assumption. We use the definition given in Vershynin (2018) (see the detailed definitions and properties of sub-Gaussian variables in Appendix B). Note that we do not require each input or disturbance to have a zero mean (see Definitions B.1 and B.4). The formal assumptions are given below.
@@
In this work, we focus on input-output analysis and aim to identify the model governing the mapping from (truncated) control inputs $(u_{t},\dots,u_{t-\tau})$ to observation outputs $y_{t}$, where $\tau$ denotes the input memory length specified by the user to construct the mapping. To this end, we rewrite (1) as follows:
$y_{t}$ $=g(x_{t},u_{t})=g(f(x_{t-1},u_{t-1},w_{t-1}),u_{t})=\cdots$
The following lemma separates the effect of disturbances and the oldest state with the effect of inputs, which allows us to establish the input-output mapping. The proof can be found in Appendix C.
Lemma 2.5.
For the equation (6), note that the term $g\circ f\circ\cdots\circ f$ is a function of $u_{t},u_{t-1},\dots,u_{t-\tau}$. We convert this nonlinear function to a linear combination of basis functions taking a truncated number of control inputs, in which we establish a time-invariant system in the sense that the number of inputs is fixed:
$g(f(\cdots f(f(0,u_{t-\tau},0),u_{t-\tau+1},0),\cdots,u_{t-1},0),u_{t})=G^{*}\cdot[\phi_{1}(\bm{U}_{t}^{(\tau)})~{}\cdots~{}\phi_{M}(\bm{U}_{t}^{(\tau)})]^{T},$ (7)
where $\bm{U}_{t}^{(\tau)}=(u_{t},\dots,u_{t-\tau})\in(\mathcal{U})^{\tau+1}$ is the stack of inputs from the time $t-\tau$ to $t$, the basis functions $\phi_{i}:(\mathbb{R}^{m})^{\tau+1}\to\mathbb{R}$ are distinct nonlinear mappings for $i=1,\dots,M$, and the matrix $G^{*}\in\mathbb{R}^{r\times M}$ explains how the nonlinear transformation of the inputs is mapped to the observation outputs. The number of basis functions $M$ as well as how to design them can be chosen at the discretion of the user. Each basis function needs to be satisfy some mild assumptions formally presented as below.
Assumption 2.6 (Basis functions).
$|\phi_{i}(\bm{U}_{t}^{(\tau)})-\phi_{i}(\bm{\tilde{U}}_{t}^{(\tau)})|\leq L_{\phi}(\|u_{t}-\tilde{u}_{t}\|_{2}+\cdots+\|u_{t-\tau}-\tilde{u}_{t-\tau}\|_{2}),\quad\forall i=1,\dots,M,$ (8)
for all $u_{t},\tilde{u}_{t},\dots,u_{t-\tau},\tilde{u}_{t-\tau}\in\mathcal{U}$. Let $\Phi(\bm{U}_{t}^{(\tau)}):=[\phi_{1}(\bm{U}_{t}^{(\tau)})~{}\cdots~{}\phi_{M}(\bm{U}_{t}^{(\tau)})]^{T}$. Also, each basis function with inputs should excite the system for the exploration in learning the system. In other words, for all $t\geq\tau$, there exists a positive constant $\lambda$ such that
$\mathbb{E}\left[\Phi(\bm{U}_{t}^{(\tau)})\Phi(\bm{U}_{t}^{(\tau)})^{T}\right]\succeq\lambda^{2}I_{M}.$ (9)
We further assume that $\Phi(0)=0$, meaning that zero input results in zero basis function values.
@@
While most of the existing literature typically assumes zero-mean Gaussian inputs, we significantly relax this assumption by only requiring $u_{t}$ to be sub-Gaussian (see Assumption 2.3) and $\Phi(\bm{U}_{t}^{(\tau)})$ to be Lipschitz and excite the system (see Assumption 2.6). A simple example that satisfies these assumptions is the aforementioned standard case of Gaussian inputs $N(0,\lambda^{2}I_{m})$. If the basis function has a simple form of $\Phi(\bm{U}_{t}^{(\tau)})=[u_{t}^{T}~{}\cdots~{}u_{t-\tau}^{T}]^{T}$, this implies that $\mathbb{E}[\Phi(\bm{U}_{t}^{(\tau)})\Phi(\bm{U}_{t}^{(\tau)})^{T}]=\lambda^{2}I_{\tau m}$, which satisfies the condition (9). Thus, Assumptions 2.3 and 2.6 characterize the conditions on control inputs in a much more general form applicable to the nonlinear system identification.
For a generic system with control inputs, if the disturbance $w_{t}$ is always adversarial with nonzero-mean, any estimator may be misled by the adversary—even in a fully observed system. We present a concrete example; a related case without control inputs appears in Kim and Lavaei (2024). Consider the scalar linear system $x_{t+1}=a^{*}x_{t}+b^{*}u_{t}+w_{t}$, where the true system is $a^{*}=0.8$, $b^{*}=0.1$, and $x_{0}=1$. Suppose that $u_{t}$ is $1$ or $-1$, each with probability $1/2$, and $w_{t}$ is chosen by an adversary as $-3x_{t}\cdot\text{sgn}(x_{t})$ for all $t\geq 0$. In such a case, one always attains
$x_{t}>0.1~{}\text{for}~{}t=0,2,4,\dots,\quad x_{t}<-0.1~{}\text{for}~{}t=1,3,5,\dots.$
However, a valid estimation method may try to minimize the distance between $x_{t+1}$ and $\hat{a}x_{t}+\hat{b}u_{t}$ to find an estimate $\hat{a},\hat{b}$. Given that $\hat{b}$ is correctly identified, one can only arrive at a negative estimate $\hat{a}$ despite the true system $a^{*}>0$, since we have $\text{sgn}(x_{t})\neq\text{sgn}(x_{t+1}-b^{*}u_{t})$ for all $t\geq 0$.
Hence, we may need to restrict the time instances $t$ in which the adversary may be able to fully attack the system via $w_{t}$. We now formally present the additional restriction on our disturbances $w_{t}$, under which the input-output mapping in (6) can accurately be estimated.

### v2

@@ Abstract
This paper is concerned with learning the input-output mapping of general nonlinear dynamical systems. While the existing literature focuses on Gaussian inputs and benign disturbances, we significantly broaden the scope of admissible control inputs and allow correlated, nonzero-mean, adversarial disturbances. With our reformulation as a linear combination of basis functions, we prove that the $\ell_{2}$-norm estimator overcomes the challenges as long as the probability that the system is under adversarial attack at a given time is smaller than a certain threshold. We provide an estimation error bound that decays with the input memory length and prove its optimality by constructing a problem instance that suffers from the same bound under adversarial attacks. Our work provides a sharp input-output analysis for a generic nonlinear and partially observed system under significantly generalized assumptions compared to existing works.
@@ 1 Introduction
Dynamical systems describe how the state of a system evolves over time according to specific laws. Such systems are ubiquitous in scientific and engineering disciplines, including computer networks (Low et al., 2002), deep learning (Meunier et al., 2022), portfolio management (Grinold & Kahn, 2000), biology (Murray, 2007), and optimal control (Dorf & Bishop, 2011). In many practical settings, however, the underlying dynamics are too complex to be explicitly characterized, resulting in models with partially or entirely unknown parameters. Designing controllers or making predictions without first identifying these unknowns can lead to suboptimal or even unsafe outcomes. To address this challenge, the field of system identification focuses on identifying system dynamics from observed input-output data.
There has been extensive research in system identification under various structural and disturbance assumptions (Simchowitz et al., 2018; Faradonbeh et al., 2018; Simchowitz et al., 2019; Jedra & Proutiere, 2020; Sarkar et al., 2021; Oymak & Ozay, 2022; Yalcin et al., 2024; Zhang et al., 2025; Kim & Lavaei, 2025a; c) While these works provide strong theoretical guarantees and practical algorithms, the majority of them concentrate on linear systems. However, many real-world systems are inherently nonlinear (Grinold & Kahn, 2000; Low et al., 2002; Murray, 2007), which motivates us to develop identification methods that go beyond the linear setting.
@@
$x_{t+1}$ $=f(x_{t},u_{t},w_{t}),~\quad~y_{t}=g(x_{t},u_{t}),\quad t=0,1,\dots,T-1,$ (1)
@@
where $x_{t}\in\mathbb{R}^{n}$ is the state, $u_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, and $y_{t}\in\mathbb{R}^{r}$ is the observation at time $t$. The set $\mathcal{U}$ consists of admissible control inputs, and $T$ is the time horizon. Under adversarially chosen disturbances $w_{t}\in\mathbb{R}^{d}$, our goal is to identify the input-output mapping of the system (1). To be specific, given an input memory length $\tau>0$, we study the mapping from the recent input sequence $(u_{t},\dots,u_{t-\tau})$ to the observation $y_{t}$.
(a) Approximation of the true input-output mapping via a linearly parameterized mapping
(b) Impact of detectable attacks versus occasional undetectable attacks on outputs
Figure 1: Input-output analysis of linearly parameterized mappings under clean and corrupted outputs. (a) We reformulate the input-output behavior of nonlinear systems as a linearly parameterized nonlinear system, which can be approximated to arbitrary precision $\epsilon$, given a sufficiently expressive set of basis functions $\phi_{1},\dots,\phi_{M}$. (b) We assume that most attacks are detectable and produce clean outputs, whereas undetectable attacks occur infrequently but can have arbitrarily large magnitudes, producing completely corrupted outputs. Our goal is to identify a linearly parameterized input-output mapping from this partially corrupted output trajectory.
In this paper, we approximate the input-output behavior of the system (1) using a finite-memory reformulation, which offers a tractable and expressive representation of the system under mild assumptions:
$\text{(schematic)}\quad y_{t}=G^{*}\cdot\Phi(u_{t},\dots,u_{t-\tau})+\textit{residual terms}+\textit{approximation error vector},$ (2)
@@
where $M$ is the number of basis functions, $\Phi:(\mathcal{U})^{\tau+1}\to\mathbb{R}^{M}$ is the stack of basis functions, $G^{*}\in\mathbb{R}^{M\times r}$ represents the matrix governing the input-output mapping, $\tau>0$ is input memory length, and the residual terms are functions of disturbances and “far” past states. Note that the far past states are exponentially small with the exponent $\tau$ under stability conditions. We will formalize this schematic form in Section 2.
In particular, given that the basis functions $\Phi(\cdot)$ are sufficiently expressive, it is guaranteed from the function approximation theory that a wide class of nonlinear mappings can be approximated to arbitrary precision, up to an approximation error vector of an arbitrarily small norm, using a finite set of appropriately chosen basis functions such as radial basis functions (RBF) (Chen et al., 1991), Volterra kernels (Boyd & Chua, 1985), and random feature models (Rahimi & Recht, 2007). Motivated by this, we focus on analyzing the corresponding linearly parameterized approximation of the input–output mapping (see Figure 1(a)).
Allowing for a small approximation error, we reduce the system identification task to estimating $G^{*}$. However, adversarial disturbances $w_{t}$, combined with the partial observability of the nonlinear system, introduce significant challenges to accurately recovering $G^{*}$. We categorize adversarial disturbances as either detectable or undetectable attacks: the former are reliably detected and corrected by a well-designed detector and feedback controller (Fawzi et al., 2014; Shoukry & Tabuada, 2016; Pajic et al., 2017), whereas the latter—though injected occasionally—corrupt the outputs and hinder the identification of the mapping $G^{*}$ (see Figure 1(b)). Furthermore, it is natural to ask whether restrictions on admissible control inputs $u_{t}$ may also impede the identification task. To this end, we pose the following central question:
When and how can we accurately estimate the true $G^{*}$ under
nonzero-mean, non-Gaussian inputs and correlated, nonzero-mean, adversarial disturbances?
@@
Dynamics
Available
Outputs
Gaussian
input
i.i.d.
Disturbance
Zero-mean
Disturbance
Identification
Approach
@@
Oymak & Ozay (2022) Linear Partial Yes Yes Yes Parametric
@@
Kim & Lavaei (2025c) Linear Partial Yes No No Parametric
@@
3) We establish that the estimation error of identifying $G^{*}$ using the $\ell_{2}$-norm estimator is $O(\rho^{\tau})$, where $0<\rho<1$ is the contraction factor of the function $f$. Notably, we further provide a matching lower bound of $\Omega(\rho^{\tau})$ on the estimation error, showing that the presented bound is indeed optimal.
Related works. We focus on identifying the input-output mapping of the system, since in many settings it suffices to capture how control actions influence observable outcomes (Abbeel et al., 2006; Deisenroth & Rasmussen, 2011). For instance, in model-based reinforcement learning (RL), the agent first learns an input–output model of the environment and subsequently uses it to make informed decisions (Moerland et al., 2023). To ensure tractability of our analysis, we adopt a parameterized system with a finite-memory approximation, which yields interpretable and computationally efficient models—particularly when the chosen function class closely aligns with the true system dynamics (Chen, 1995; Giannakis & Serpedin, 2001). The finite-memory approach is consistent with classical nonlinear system identification methods, such as Volterra series truncations (Boyd & Chua, 1985) and NARMAX models (Billings, 2013). Further details on related works are in Appendix A.
@@
Outline. The paper is organized as follows. In Section 2, we formulate the problem and state the relevant assumptions. In Section 3, we prove that the $\ell_{2}$-norm estimator achieves the optimal estimation error and provides the analysis outline. In Section 4, we present numerical experiments to validate our main results. Finally, concluding remarks are provided in Section 5.
@@
Notation. Let $\mathbb{R}^{n}$ denote the set of $n$-dimensional vectors and $\mathbb{R}^{n\times n}$ denote the set of $n\times n$ matrices. For a matrix $A$, $\|A\|_{F}$ denotes the Frobenius norm of the matrix. For a vector $x$, $\|x\|_{2}$ denotes the $\ell_{2}$-norm of the vector. For a set $S$, the $k$-fold Cartesian product $S\times S\times\cdots\times S$ (with $k$ factors) is denoted by $(S)^{k}$. For an event $E$, the indicator function $\mathbb{I}\{E\}$ equals $1$ if $E$ occurs, and 0 otherwise. $\mathbb{P}(E)$ denotes the probability that the event occurs. We use $O(\cdot)$ for the big-$O$ notation and $\Omega(\cdot)$ for the big-$\Omega$ notation. Let $I_{n}$ denote the $n\times n$ identity matrix. The notation $\succeq$ denotes positive semidefiniteness. Let $N(\mu,\Sigma)$ denote the Gaussian distribution with mean $\mu$ and covariance $\Sigma$, and $\text{Unif}[a,b]^{n}$ denote the uniform distribution on the hypercube $a,b]^{n}\subset\mathbb{R}^{n}$. Finally, let $\mathbb{E}$ denote the expectation operator.
@@ 2 Problem Formulation
In (1), we study a nonlinear dynamical system $x_{t+1}=f(x_{t},u_{t},w_{t})$ and $y_{t}=g(x_{t},u_{t})$, where the state equation is governed by the dynamics $f:\mathbb{R}^{n}\times\mathcal{U}\times\mathbb{R}^{d}\to\mathbb{R}^{n}$ and the observation equation is determined by the measurement model $g:\mathbb{R}^{n}\times\mathcal{U}\to\mathbb{R}^{r}$. We have the discretion to design control inputs $u_{0},u_{1},\dots,u_{T-1}$ and we have access to a single observation trajectory consisting of partial observations $y_{0},y_{1},\dots,y_{T-1}$. We assume the Lipschitz continuity of the measurement model $g$ and the contraction property for the dynamics $f$ to ensure system stability and prevent the explosion of the nonlinear system, which is common in control theory literature (Tsukamoto et al., 2021; Lin et al., 2023). The formal assumption on the dynamics is given below.
@@
$g$ is Lipschitz continuous; i.e., there exists $L>0$ such that
@@
for all $x,\tilde{x}\in\mathbb{R}^{n}$, $u,\tilde{u}\in\mathcal{U}$. Moreover, note that for $k\geq 1$, the $k$-fold composition of the dynamics $f$, denoted by $f^{(k)}$, maps $(x_{t-k},u_{t-k},\dots,u_{t-1},w_{t-k},\dots,w_{t-1})$ to $x_{t}$. We assume that $f^{(k)}$ is Lipschitz continuous in its oldest arguments $(x_{t-k},u_{t-k},w_{t-k})$ with constant $C\rho^{k}$ for some $C>0$ and $0<\rho<1$, with later inputs and disturbances fixed. In other words, we have
@@
In this work, we focus on input-output analysis and aim to identify the model governing the mapping from (truncated) control inputs $(u_{t},\dots,u_{t-\tau})$ to observation outputs $y_{t}$, where $\tau$ denotes the input memory length specified by the user to construct the mapping. As described in the introduction, we will reformulate the true mapping to a linearly parameterized input-output mapping with a finite-memory approximation. To this end, we outline the following four steps.
Step 1 By recursively applying the system dynamics (1), the observation $y_{t}$ can be represented as
$y_{t}$ $=g(x_{t},u_{t})=g(f(x_{t-1},u_{t-1},w_{t-1}),u_{t})=\cdots$
Step 2 We next separate the effect of disturbances and the oldest state with the effect of inputs, which allows us to establish the input-output mapping. The result is summarized in the following lemma (see the proof in Appendix C).
Lemma 2.3.
Step 3 For the equation (6), note that the term $g\circ f\circ\cdots\circ f$ is a function of $u_{t},u_{t-1},\dots,u_{t-\tau}$. Allowing for a small approximation error vector $\bm{\epsilon}_{t}$, we convert this nonlinear function to a linear combination of basis functions taking a truncated number of control inputs, in which we establish a time-invariant system in the sense that the input memory length is fixed:
$g(f(\cdots f(f(0,u_{t-\tau},0),u_{t-\tau+1},0),\cdots,u_{t-1},0),u_{t})=G^{*}\cdot[\phi_{1}(\bm{U}_{t}^{(\tau)})~\cdots~\phi_{M}(\bm{U}_{t}^{(\tau)})]^{T}+\bm{\epsilon}_{t},$ (7)
where $\bm{U}_{t}^{(\tau)}=(u_{t},\dots,u_{t-\tau})\in(\mathcal{U})^{\tau+1}$ is the stack of inputs from the time $t-\tau$ to $t$, the basis functions $\phi_{i}:(\mathbb{R}^{m})^{\tau+1}\to\mathbb{R}$ are distinct nonlinear mappings for $i=1,\dots,M$, and the matrix $G^{*}\in\mathbb{R}^{r\times M}$ explains how the nonlinear transformation of the inputs is mapped to the observation outputs. The number of basis functions $M$ as well as how to design them can be chosen at the discretion of the user.
Step 4 Let $\Phi(\bm{U}_{t}^{(\tau)}):=[\phi_{1}(\bm{U}_{t}^{(\tau)})~\cdots~\phi_{M}(\bm{U}_{t}^{(\tau)})]^{T}$. Considering the relationships (6) and (7), we finally arrive at the equation
$y_{t}=G^{*}\cdot\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}$ (8)
for all $t\geq\tau$. This provides an equivalent representation of $y_{t}$ via a linearly parameterized mapping (see Figure 1(a)), with the goal of accurately estimating the true matrix $G^{*}$ that governs the input-output mapping from the control inputs $(u_{t},\dots,u_{t-\tau})$ to the observation $y_{t}$. ∎
We now present the mild assumptions that each basis function should satisfy.
Assumption 2.4 (Basis functions).
$|\phi_{i}(\bm{U}_{t}^{(\tau)})-\phi_{i}(\bm{\tilde{U}}_{t}^{(\tau)})|\leq L_{\phi}\|\bm{U}_{t}^{(\tau)}-\bm{\tilde{U}}_{t}^{(\tau)}\|_{2},\quad\forall i=1,\dots,M,$ (9)
for all $\bm{U}_{t}^{(\tau)},\bm{\tilde{U}}_{t}^{(\tau)}\in(\mathcal{U})^{\tau+1}$. Also, each basis function with inputs should excite the system for the exploration in learning the system. In other words, there exists a universal constant $\lambda>0$ such that
$\mathbb{E}\left[\Phi(\bm{U}_{t}^{(\tau)})\Phi(\bm{U}_{t}^{(\tau)})^{T}\right]\succeq\lambda^{2}I_{M}$ (10)
holds for all $t\geq\tau$. We further assume that $\Phi(0)=0$, meaning that zero input results in zero basis function values.
We also consider both inputs and disturbances on the system to be sub-Gaussian variables. For example, any bounded variables automatically satisfy this assumption. We use the definition given in Vershynin (2018) (see the detailed definitions and properties of sub-Gaussian variables in Appendix B). Note that we do not require each input or disturbance to have a zero mean (see Definitions B.1 and B.4). The formal assumptions are given below.
@@
While existing literature typically assumes zero-mean Gaussian inputs, we significantly relax this assumption by only requiring $u_{t}$ to be sub-Gaussian (see Assumption 2.5) and $\Phi(\bm{U}_{t}^{(\tau)})$ to be Lipschitz and excite the system (see Assumption 2.4), without imposing a zero-mean condition. These assumptions characterize the general conditions on control inputs applicable to nonlinear system identification. In practice, control inputs are often synthesized in the form of $K(y_{t})+[\text{sub-Gaussian excitation term}]$, where $K$ is a controller and $y_{t}$ is the observation at time $t$. With excitation terms, $K(y_{t})$ may indeed have a nonzero mean, which incorporates realistic scenarios.
If the disturbance $w_{t}$ is always adversarial with nonzero-mean, any estimator may be misled. For example, the adversary can always inject an attack $w_{t}$ that drives the next state $x_{t+1}$ to be irrelevant to the current state $x_{t}$, preventing any valid estimation method from extracting useful information (Kim & Lavaei, 2025a). Hence, we may need to restrict the time instances $t$ in which the adversary may be able to fully attack the system via $w_{t}$. We now formally present the additional restriction on our disturbances $w_{t}$, under which the input-output mapping in (8) is accurately estimated.